### v1

# Reflection on Data Storytelling Tools in the Generative AI Era from the Human-AI Collaboration Perspective

Abstract.
Human-AI collaborative tools attract attentions from the data storytelling community to lower the barrier of expertise and streamline the workflow. The recent advance in large-scale generative AI techniques, e.g., large language models (LLMs) and text-to-image models, has the potential to enhance data storytelling with their power in visual and narration generation. After two years since these techniques were publicly available, it is important to reflect our progress of applying them and have an outlook for future opportunities. To achieve the goal, we compare the collaboration patterns of the latest tools with those of earlier ones using a dedicated framework for understanding human-AI collaboration in data storytelling. Through comparison, we identify persistent collaboration patterns, e.g., human-creator + AI-assistant, and emerging ones, e.g., AI-creator + human-reviewer. The benefits of these AI techniques and other implications to human-AI collaboration are also revealed. We further propose future directions to hopefully ignite innovations.
A browser of our collected tools is available at https://human-ai.notion.site/data-storytelling-dec-2024.
Data storytelling, human-AI collaboration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Visualization
1. Introduction
Data storytelling is considered as one of the major research directions in visualization research (Kosara and Mackinlay, 2013). To lower the barrier of telling appealing and effective stories, researchers have spent considerable efforts to build AI-powered tools to facilitate their creation and communication with different strategies of human-AI collaboration (Li et al., 2024b). In these tools, AI collaborators are often powered by heuristic-based methods (Wang et al., 2019), traditional machine learning models (Fu et al., 2019), or smaller-scale deep learning models (Qian et al., 2020).
Compared to the previous techniques for AI collaborators, the recently emerging large-scale generative AI models, including the text-to-image models (Yang et al., 2023) and large language models (LLMs) (Zhao et al., 2023), can achieve better performance on various data storytelling-related tasks, such as data analysis (Inala et al., 2024) and text generation (Zheng et al., 2023), and enhance the communication between humans and AI with conversations. These advantages indicate their potentials to be game-changers in the research direction of human-AI collaboration for data storytelling, including improving the experience of collaborating with AI and diversifying the collaboration patterns between humans and AI (Li et al., 2024b). After two years of the public access of these models, it is a critical time point to reflect how this research discipline progresses in the new era of large-scale generative AI models and identify future opportunities. To achieve the goal, it is essential not only to focus on how these generative AI techniques are applied in existing tools, as explored in a previous survey (He et al., 2024), but more importantly, to compare the human-AI collaboration patterns in the latest tools in the generative AI era with those in earlier ones. Only through this comparison can we understand the shift in human-AI collaboration paradigms, identify the value of these powerful techniques in enhancing human-AI collaboration, and propose future research directions.
To perform the comparison, the foundation is a fine-grained and technique-agnostic framework to delineate human-AI collaboration patterns in the tools both before and in the era of generative AI. We notice that the framework by Li et al. (Li et al., 2024b) can fulfill the requirement. The framework breaks down the human-AI collaboration in data storytelling tools to two dimensions, the stages in the workflow (e.g., planning and implementation) and the roles of human and AI collaborators (e.g., creator and optimizer), which are both disentangled from the AI techniques applied in tools. The authors further leveraged the framework to conduct a comprehensive survey of related tools before mid-2023. Based on the framework and previously surveyed tools, our work reviews the latest 27 data storytelling tools and compares their human-AI collaboration patterns with those of earlier tools before the public availability of these large-scale AI models.
Through the comparison, we find patterns that remain popular in the new era, such as the mode where AI assists human creators, and emerging patterns, such as human as reviewers to AI-created content. Then, we discuss their implications to human-AI collaboration, such as the advantages of these models, the necessity and suitable scenarios of both human-AI collaborative and fully automatic tools, and noteworthy research opportunities, such as the dependency of roles in different stages. By reflecting on the progress so far and exploring future directions, we hope this paper serves as a cornerstone for investigating the collaboration between humans and large-scale generative AI.
2. Related Surveys
To review data storytelling tools, there were several previous surveys for different focuses. The first group of surveys (Tong et al., 2018; Zhao and Elmqvist, 2023; Ren et al., 2023) reviews tools from the aspect of how they support narrative techniques, such as data story structures (Tong et al., 2018) and components (Zhao and Elmqvist, 2023; Shen et al., 2025). The second group focuses on the application of automation techniques, such as AI, in data storytelling tools. Schröder et al. (Schröder et al., 2023) only grouped AI-powered automatic tools based on their functions and genres. Chen et al. (Chen et al., 2023a) further categorized AI-powered tools into manual, mixed-initiative, and automatic types. Li et al. (Li et al., 2024b) proposed a more fine-grained framework to explicitly understand the collaboration mechanism between human and AI through their collaboration stages and roles. Recently, motivated by the advancement of generative AI models, He et al. (He et al., 2024) reviewed their usage in tools based on tasks, such as insight identification and narration generation.
Under the second category, our paper aims to identify the insights about how human-AI collaboration in data storytelling tools is advanced in the era of generative AI models. It serves as a sequel to the work by Li et al. (Li et al., 2024b) to refresh our understanding through comparing the human-AI collaboration patterns of the latest tools with the patterns of earlier tools in their work. Moreover, unlike the survey by He et al. (He et al., 2024) and others at the intersection of large-scale generative AI models and visualization (Ye et al., 2024; Yang et al., 2024), we focus less on the technical perspective of applying these models but emphasize the advancement of higher-level human-AI collaboration patterns of data storytelling tools.
3. Method and Overview of Tools
Following Li et al. (Li et al., 2024b), we review the recent data storytelling tools after June 2023 and compare their collaboration patterns with those of earlier ones. This section introduces our detailed method and an overview of tools. For brevity, we denote the tools after June 2023 as the latest tools in the era of generative AI and the tools before as earlier tools.
3.1. Tool Collection
In our paper, we focus on the tools in the academia since their designs are often described with more details. To collect the corpus, we searched tools between June 2023 and Dec. 2024 in related premier human-computer interaction venues, including CHI, UIST, and IUI, and visualization venues, including VIS, PacificVis, EuroVis, and TVCG, with keywords “narrative visualization” and “data storytelling”. Finally, we collected 27 tools, as shown in Table 1.
After a careful consideration, we decided to compare human-AI collaboration patterns of the latest tools after June 2023 with those of earlier tools to study the research trend of human-AI collaborative tool in the era of generative AI. First, the dates of public access to two representative large-scale generative AI techniques, ChatGPT and Stable Diffusion, are Nov. 30, 2022 (OpenAI, 2022) and Aug. 22, 2022 (Stability.ai, 2023), respectively. Considering the delay led by tool development and paper publishing, the research tools to apply them and other similar techniques might only be available in mid-2023. Then, we examined the corpus of Li et al. (Li et al., 2024b), which includes tools published before June 2023. Their corpus do not have extensive application of these large-scale generative models, while the authors have discussed the potentials to apply them. It also supports our decision. Our surveyed tools further verify the decision: there were pioneering ones that apply text-to-image models (Xiao et al., 2023) and LLMs (Sultanum and Srinivasan, 2023; Shen et al., 2023; Lin et al., 2023) published in the VIS 2023 conference in Oct. 2023.
Table 1. This table presents the human-AI collaboration patterns of our collected tools with the framework of stages and roles by Li et al. (Li et al., 2024b) and the applied generative AI techniques. The collaborators’ roles are shown in abbreviation: H-C, H-A, H-O, H-R are the creator, assistant, optimizer, and reviewer roles performed by humans; A-C, A-A, A-O, A-R are the corresponding roles by AI. The techniques are denoted with the icons: * refers to LLMs, such as GPT-4 (OpenAI, 2023) and # refers to text-to-image models, such as diffusion models (Yang et al., 2023). The tools are sorted according to their covered stages and collaboration patterns.
[…]
This table includes six columns: the publication years of tools, the tool name, and the collaboration patterns in four stages: analysis, planning, implementation, and communication. There are in total 27 rows, where each row records one tool. The tools are sorted by their covered stages and collaboration patterns. Year Tool Analysis Planning Implementation Communication 2023 SwimFlow (Yao et al., 2023) N/A N/A H-C N/A 2024 Epigraphics (Zhou et al., 2024a) N/A N/A H-C A-A ∗# N/A 2024 AnalogyMate (Chen et al., 2024) N/A N/A H-C A-A ∗# N/A 2024 Emoji Encoder (Brehmer et al., 2024) N/A N/A H-C A-A N/A 2024 CAST+ (Shen et al., 2024c) N/A N/A H-C A-A N/A 2024 Cai et al. (Cai et al., 2024) N/A N/A H-C A-A N/A 2023 DataTales (Sultanum and Srinivasan, 2023) N/A N/A A-C H-O N/A 2024 VisTellAR (Tong et al., 2024) N/A H-C H-C A-A N/A 2024 WonderFlow (Wang et al., 2024b) N/A H-C H-C A-A N/A 2024 VisConductor (Femi-Gege et al., 2024) N/A H-C H-C H-C A-A 2023 Data Player (Shen et al., 2023) N/A H-C A-C N/A 2024 Data Playwright (Shen et al., 2024b) N/A H-C A-C H-O N/A 2023 ChartSpark (Xiao et al., 2023) H-C A-A N/A H-C A-A # N/A 2023 InkSight (Lin et al., 2023) H-C N/A A-C H-O N/A 2024 OutlineSpark (Wang et al., 2024a) H-C H-C A-A A-C H-O N/A 2024 LEVA (Zhao et al., 2024) H-C A-A H-C A-C N/A 2024 CoInsight (Li et al., 2024a) H-C A-A H-C A-A A-C N/A 2024 Data Director (Shen et al., 2024a) A-C A-C A-C N/A 2024 Shi et al. (Shi et al., 2024) A-C A-C A-C N/A 2024 Live Charts (Ying et al., 2024) A-C A-C A-C A-C 2024 Sportify (Lee et al., 2024) A-C A-C A-C A-C 2024 AiCommentator (Andrews et al., 2024) A-C A-C A-C A-C 2023 Calliope-Net (Chen et al., 2023b) A-C H-O A-C H-O A-C H-O N/A 2024 DG Comics (Kim et al., 2024) A-C H-O A-C H-O A-C H-O N/A 2024 SNIL (Cheng et al., 2024) A-C H-O A-C H-O A-C H-O N/A 2023 Socrates (Li et al., 2024a) A-C H-R A-C H-R A-C N/A 2024 Han and Isaacs (Han and Isaacs, 2024) N/A N/A N/A H-C A-A
(a) Analysis
(b) Planning
(c) Implementation
(d) Communication
Figure 1. The change of human-AI collaboration pattern frequencies between the latest and earlier tools in four stages. In a subfigure, the color of blocks and numbers in the heatmap encodes the change of pattern frequencies in the stage. A pattern is represented by a human role on the horizontal axis and an AI role on the vertical axis. Green and positive numbers mean that the pattern is more frequent in the latest tools while red and negative numbers indicate that the pattern is less frequent. A number in brown means that the pattern is newly identified in the latest tools. An empty white cell means that the pattern has not emerged in any tools. Two bar charts on the right and the top of the heatmap show the frequency change of one role between the latest and earlier tools.
[…]
This figure shows the change of collaboration patterns in four stages. The four subfigures show the pattern changes in four stages. In the analysis stage, the pattern of human-creator only decreases 23%, human-creator and AI-assistant increases 3%, AI-creator and human-assistant decreases 6%, AI-creator and human-optimizer decreases 4%, AI-creator and human-reviewer, a new pattern, increases 7%, and AI-creator only increases 23%. In the planning stage, the pattern of human-creator only decreases 16%, human-creator and AI-assistant increases 1%, human-creator and AI-optimizer decreases 7%, AI-creator and human-optimizer decreases 10%, AI-creator and human-reviewer, a new pattern, increases 6%, and AI-creator only increases 26%. In the implementation stage, the pattern of human-creator only decreases 15%, human-creator and AI-assistant increases 10%, AI-creator and human-assistant decreases 5%, AI-creator and human-optimizer decreases 8%, AI-assistant only decreases 2%, AI-reviewer only decreases 3%, and AI-creator only increases 23%. In the communication stage, the pattern of human-creator only decreases 60%, human-creator and AI-assistant decrease 7%, and AI-creator only increases 67%. The other patterns that are not mentioned above have not been implemented in existing tools.
3.2. Pattern Coding
After collecting all tools, we coded their human-AI collaboration patterns with the framework proposed by Li et al. (Li et al., 2024b). The framework has two dimensions: the roles of human and AI collaborators (i.e., human-creator and AI-creator to complete the majority of tasks in each stage, human-assistant and AI-assistant to help creators on designated tasks in a stage, human-optimizer and AI-optimizer to improve the quality of the output of one task or the entire stage autonomously, and human-reviewer and AI-reviewer to provide suggestions for the output from a task or stage) and their collaboration stages (i.e., analysis stage to derive data insights and knowledge from datasets, planning stage to decide how to convey analytical findings through a data story, implementation stage to create data story content following the plan, and communication stage to share data stories directly with audiences). One author coded all tools after understanding the coding principles and the coded tools by Li et al. comprehensively. During the coding process, the results were also calibrated with previously coded tools as references. The final results are presented in Table 1 and an online browser: https://human-ai.notion.site/data-storytelling-dec-2024. To facilitate comparison, we computed and visualized the changes in the relative frequencies [^1] All frequencies refer to relative frequencies in this paper to accommodate the different numbers of tools in each stage in earlier and the latest tools. of collaboration patterns in Figure 1(d). The frequency of a pattern $p$ in a stage $s$ is computed as $f=Occurrence_{p\textrm{ in }s}/NumTool_{s}$ and the frequency change between the latest tools and earlier tools is $f_{latest}-f_{earlier}$. Therefore, a positive change in Figure 1(d) means that the pattern is gaining popular while a negative one shows the frequency is declining. Figures 3(d) and 4(d) show the frequencies of earlier and the latest tools.
We decided to apply the framework for a series of reasons. First, the framework provides a more fine-grained approach to distinguish the collaboration patterns of human and AI in the storytelling workflow than others (e.g.(Chen et al., 2023a)). Furthermore, the framework was summarized based on an empirical interview study (Li et al., 2023a) that envisions the potential usage and challenges of AI collaborators after these large-scale AI models were released. Checking the alignment between existing tools and user expectations can help us reflect and identify opportunities. Another advantage is that the framework is technique-agnostic, which allows us to compare the collaboration patterns without interference led by the differences in techniques. Lastly, earlier tools examined with the framework by Li et al. (Li et al., 2024b) can be easily re-used for comparison. The results were verified through peer review and thus provide solid foundation for our research.
3.3. Overview of Tools
Combining the results of ours and Li et al. (Li et al., 2024b), we notice a considerable growth in the number of data storytelling tools in 2024 (Figure 2). Compared to the 11 tools published in 2023, the 20 tools in 2024 nearly double the previous number, reaching a historic high. Furthermore, 16 out of the 27 latest tools apply generative AI models, where 13 tools use LLMs merely, one tool uses text-to-image models merely, and two tools use both (Table 1). These models are most frequently used in the implementation stage (15/27), following up by the analysis (5/27), the planning (4/27), and the communication stages (3/27). The findings reveal that large-scale generative AI models has been widely applied across the whole storytelling workflow, reinforcing our motivation of understanding changes in the generative AI era.
4. Findings
Based on Figure 1(d), this section introduces insights about the change of collaboration patterns, including the continuously popular patterns (Section 4.1), the new patterns (Section 4.2), and the patterns that attract more attention (Section 4.3).
4.1. What pattern remains popular?
The most popular patterns with both humans’ and AI’s participation in earlier tools are human-creator + AI-assistant  [^2] The plus sign indicates the two roles collaborate with each other in a specific stage. and AI-creator + human-optimizer since they allow humans to control the content while lowering the barrier of creation through automation (Li et al., 2024b). In the latest tools, their frequency only has limited fluctuation, suggesting that they are still welcomed. We noticed two noteworthy advancements brought by generative AI models when applying the patterns.
First, in the communication stage, we identify an innovative usage of AI-assistant, which may lead to future research opportunities. The usage of AI-assistant in the communication stage was to support individual users in live presentations, such as retrieving and presenting charts based on sketch (Lee et al., 2013) and supporting interaction with charts with gesture recognition (Hall et al., 2022). Building on LLMs’ ability to understand and generate texts, Han and Isaacs (Han and Isaacs, 2024) proposed an AI-assistant to record the communication process between human-creators around charts and create an interactive data document with texts, visualizations, and annotations from the conversation. The work has more implications beyond identifying a new function of AI-assistants powered by generative AI models. It demonstrates a new collaborative data story creation mode, where humans freely express their insights around data in conversations and AI documents and organizes humans’ ideas into a more structured data story. It is interesting to examine the benefit of this collaboration mode and explore other potentials. For example, will it be possible to ask AI to interview humans and write data stories or generate engaging data videos like talk shows based on humans’ conversations around data? Furthermore, from the human-AI collaboration perspective, this AI-assistant has a notable difference with others. It serves as a third-party collaborator to document multiple humans’ communication, while not collaborating with individual humans and contributing to their creation or presentation (e.g., AI-assistants in VizFlow (Sultanum et al., 2021) and SketchStory (Lee et al., 2013)). We should further investigate the potentials of human-AI collaboration to support multiple humans’ work in data storytelling (Chevalier et al., 2018).
Second, the application of large-scale generative AI can improve the collaboration experience from several perspectives. The first perspective is to increase AI collaborators’ output space significantly. In our corpus, we notice that Epigrahics (Zhou et al., 2024a), AnalogyMate (Chen et al., 2024), and ChartSpark (Xiao et al., 2023) follow a similar design in the implementation stage. Their AI-assistants apply generative models to create visual components and provide design inspirations to support human-creators for authoring static or animated infographics. The similar design exists in earlier tools, such as Vistylist (Shi et al., 2022), Chartreuse (Cui et al., 2021) and DataQuilt (Zhang et al., 2020). However, they can only support re-using existing visual elements, greatly limiting the potential space of creation. The AI collaborators powered by retrieval-based algorithms also suffer from the issue , such as AI-creators in Retrieve-then-Adapt (Qian et al., 2020) and Text-to-Viz (Cui et al., 2019). Compared to these earlier work, the latest tools powered by generative AI models are more flexible to create new visual elements beyond existing examples. Additionally, the interaction between collaborators can be more natural and closer to humans’ habits. Besides direct text instructions (Xiao et al., 2023; Zhao et al., 2024), OutlineSpark (Wang et al., 2024a) allows humans to input a structured outline which guides AI-creator to select cells from notebooks and generate structured slides. In Data Playwright (Shen et al., 2024b), users write annotated narrations to specify expected animation effect and timing along with narrations. In both tools, users’ intent are conveniently expressed to enhance their agency, and their efforts are reduced through more AI automation. The last advantage is output quality improvement. For example, LLMs are applied to enhance template-based text generation (Lin et al., 2023; Kim et al., 2024).
4.2. What pattern has newly appeared?
We are happy to see that there are new collaboration pattern appearing in the latest tools: the usage of human-reviewers in the analysis and the planning stages and AI-creators in the communication stage.
The collaboration pattern of AI-creator + human-reviewer has been discussed as a potential strategy to enhance AI automation and human agency (Li et al., 2024b). Compared to human-optimizers, human-reviewers can control the data story with feedback on AI-created content and thus save the effort to revise the content manually. In our surveyed tools, Socrates (Wu et al., 2023) allows human-reviewer to provide feedback to the direction of data analysis (e.g., what attributes in datasets to be included in the data story) and the narrative structure (e.g., highlight a data fact using the strategy of contrasting it with another one). Specifically, the AI-creator, powered by a heuristic-based algorithm, will generate questions and ask for human-reviewer to rate, compare, or select options. Then, the AI-creator can leverage the results to create a data story that aligns with the intent of human-reviewer. Though they do not explicitly compare the efforts spent with their tool and another with human-optimizer, i.e., Calliope (Shi et al., 2020), their results show that the created stories with Socrates match the intent better, which also implies that humans might save efforts for further adjustment. The study results preliminarily verify the advantages of human-reviewer to achieve more AI automation and more human agency than human-optimizer, when collaborating with an AI-creator. In addition to the authors’ suggestion of exploring free-form feedback, future research can further explore the design and application of human-reviewers, such as understanding the pros and cons explicitly and providing feedback to visual channels in data stories.
Another new pattern is the application of AI-creators for communicating data stories automatically. As revealed in a previous interview study (Li et al., 2023a), data workers often prefer not to ask AI to perform the creator role in the communication since they may diminish the underlying value of human-human communication, such as building trust. Therefore, we are particularly interested in the design of these AI-creators in the latest tools. Among the four AI-creators, Sportify (Lee et al., 2024) and AiCommentator (Andrews et al., 2024) leverage AI-creators to communicate automatically identified data insights in sports games to humans. Ying et al. (Ying et al., 2024) and Shi et al. (Shi et al., 2024) support chart and dataset understanding with AI-creators who communicate insights to humans with animations and narrations. Comparing these cases with the interview study, we can notice that the AI-creator communicates AI-created stories in these tools rather than presenting human-created stories as their proxies. It implies whether AI-creators can be accepted by humans in the communication stage may depend on AI’s roles in the previous stages. If the story is fully created by the AI collaborator in previous stages and reflect little human opinions, the role of AI-creators is more acceptable. On the other hand, when the story serves for the purpose of human-human communication and carries intensive information from human storytellers, the usage of AI-creators as a human proxy might raise concerns. Beyond its implication to AI-creators, the observation also has reveals that the application of different roles across stages can be dependent on each other. It can be interesting to further understand the factors that lead to the dependency and explore whether the dependency of roles exists in other domains.
4.3. What pattern is getting more popular?
Across all four stages, we notice the growing interest of the AI-creator-only pattern, where various fully automatic tools across multiple stages emerge (Ying et al., 2024; Lee et al., 2024; Shen et al., 2024a; Andrews et al., 2024). We consider that the observation is led by the capability of large-scale generative models, such as the power of analyzing data (Inala et al., 2024) and generating corresponding text explanation (Dibia, 2023). One might question if the trend contradicts with the argument that human-AI collaboration in data storytelling would gain more interests (Li et al., 2024b). Furthermore, does the advancement of AI capability mean that we should try to automate everything with AI in data storytelling and thus the research on human-AI collaboration is no longer needed?
Inspired by the discussion where collaboration patterns can be affected by tool usage scenarios (Li et al., 2024b), we dive into the scenarios of these fully automatic tools to figure out why they are designed to be AI-creators-only. In the usage scenarios of Sportify (Lee et al., 2024) and AiCommentator (Andrews et al., 2024), sports fans issue targeted queries about matches for desired information, such as tactics or player performances. Then, AI collaborators create tailored and casual data stories based on the queries and communicate them to humans automatically. Compared to other scenarios, this one emphasizes the need of being customized and real-time while tolerating that the quality might not be perfect due to the casual nature of the scenario. When telling data stories to the public (Segel and Heer, 2010), the goal is often set by the authors and thus the stories are not tailored for a reader. In the scenario of Sportify and AiCommentator, data stories are told based on an individual user’s intent, implying that they have to be customized. Furthermore, making serious data stories for the public requires considerable time to polish them (Chevalier et al., 2018), while real-time processing might be prioritized in casual and low-stake scenarios (Lee et al., 2024; Andrews et al., 2024).
To fulfill the requirements of customization and high responsiveness, the creation and communication of data stories can hardly be intervened by human-creators or human-assistants, who are likely to take longer time and higher financial cost to finish tasks (Li et al., 2023a). The traditional AI algorithms may also have limited flexibility for those highly customized data stories. The emergence of the latest large-scale AI models can fill the gap due to their ability of consuming free-form input to describe audiences’ goals and generating and communicating data stories (Section 4.1). They can serve as AI-creators to automate the whole workflow. Furthermore, since the scenarios are casual and low-stake, careful examination and quality improvement by human-optimizers or human-reviewers might not be needed.
The scenario of Live Charts (Ying et al., 2024) and the work by Shi et al. (Shi et al., 2024) impose similar requirements to storytellers. The authors propose to animate static charts and support them with narration to facilitate understanding. The generated data stories should help user to interpret promptly but their designs do not need to be polished. Considering the high cost of humans further, a fully automatic pipeline is more suitable. Another similar audience-oriented scenario is to enhance text-only data story understanding by generating visuals (Li et al., 2024b), where automatic tools (Hullman et al., 2013; Gao et al., 2014; Shao et al., 2025) are designed.
After examining these cases, we can attempt to answer the question about the necessity and potential future research direction of human-AI collaboration. First, led by the outstanding capability of large-scale generative AI models, it is a natural and irresistible trend that the application of AI-creators will be wider. To cope with the trend, we should actively experiment with these models (Li et al., 2024b) and learn how to apply them smartly, such as effective prompt design (Shen et al., 2024a). Furthermore, we should realize that the application of AI-creators is not only related to their performance but the scenario as well. The scenarios of fully automatic tools share common features, such as low-stake scenarios that require real-time generation of personalized stories. Additionally, there can be considerations regarding human-human communication (Section 4.2). In other scenarios, fully automatic tools might not be the suitable choice. For example, in contrast to the aforementioned tools (Lee et al., 2024; Andrews et al., 2024), SNIL (Cheng et al., 2024) is designed for the scenario where data journalists create serious content for the public. As a result, though SNIL can generate data articles for basketball games, it emphasizes the collaboration between AI-creators and data journalists as human-optimizers. Furthermore, in scenarios requiring creativity, the research of human-AI collaborative tools is also booming (e.g.(Chen et al., 2024; Xiao et al., 2023; Zhou et al., 2024a)). More research is needed to understand the relationship between the needs for different AI collaborators and the characteristics of scenarios. Then, in those scenarios requiring human-AI collaboration, designing suitable collaboration patterns should be studied. Finally, though these tools appears lack of human-AI collaboration from the data storytelling perspective, they support human-AI collaboration from a broader angle. For example, AI-created data stories help sports fans to understand tactics (Lee et al., 2024) and assist novices to comprehend charts (Ying et al., 2024). To conclude, we believe that human-AI collaboration remains an important topic to investigate and there will be even more opportunities to apply data storytelling for broader human-AI collaboration in data-related tasks.
5. Discussion
Section 4 presents exciting research progress in the era of generative AI, including new human-AI collaboration patterns and interactions in data storytelling tools. Meanwhile, we realize that it is not the end of this direction. There are plenty of needs and opportunities to continue our research, e.g., understanding these new patterns and investigating under-explored patterns (Figure 1(d)). This section provides suggestions for advancing this field with powerful AI models.
Large-scale Generative AI Models as Universal Engines with Unified Interfaces. In addition to the benefits in Section 4.1, a more valuable characteristic of these generative AI models, especially LLMs, is to provide universal engines and interfaces for various roles and stages. Earlier tools often equip an AI collaborator with a specialized algorithm in one stage, e.g., fact mining and organization (Wang et al., 2019; Li et al., 2023b). Next, an interface, or “shared representation” (Heer, 2019) is required to support the communication among AI collaborators in different stages and even between AI and humans (Li et al., 2024b), e.g., chart or insight specification (Tang et al., 2017; Kim and Heer, 2021). Existing research shows that LLMs can support these stages as different roles with appropriate prompt design (Lee et al., 2024; Zhao et al., 2024). Furthermore, they can leverage natural languages as a unified interface for communication, with supplementary ways of programming languages. Therefore, these models can reduce tool development cost led by specialized algorithms and release the constraint of communication channels. Researchers can apply them to investigate expected but overlooked patterns, e.g., AI-optimizer + human-optimizer (Li et al., 2024b) and human-creator + AI-reviewer (Li et al., 2023a), conveniently. Furthermore, it will be exciting to explore fluid human-AI collaboration with user-configured patterns.
Guidelines for Applying AI Models. Though powerful, these large-scale generative AI models have downsides, such as long response time (Li et al., 2025), and even risks like hallucination (Lee et al., 2023). As a result, their usage requires more research at a meta level to summarize when, where, and how they should be leveraged to achieve more benefit while mitigating their issues in data storytelling. For example, it is necessary to consider how to decompose tasks and design prompts (Shen et al., 2024a) and balance its usage with traditional algorithms to enhance the processing speed, accuracy, reliability, and etc. An design trade-off discussed by Socrate authors (Wu et al., 2023) is the usage of LLMs to extract intent for free-form feedback input by human-reviewers or only supporting limited input types but receiving more direct preferences. It will be interesting to propose AI application guidelines based on techniques, usage scenarios (Section 4.3), and collaboration patterns.
Necessity for Continuous Reflection. In the generative AI era, we are witnessing a fast growth of data storytelling tools (Section 3.3). The research progress should be reflected continuously and more regularly to summarize lessons and identify opportunities. For example, Li et al. only discussed the existence of the relationship between collaboration patterns and usage scenarios with an example of AI-creators to support data story understanding (Li et al., 2024b). Powered by generative AI, more AI-creator-only scenarios emerges, allowing us to summarize their common characteristics, such as the casual and audience-oriented nature (Section 4.3). These findings greatly enrich our knowledge about the effect of usage scenarios on tool design. The example advises that human-AI collaboration evolves with technical advancements. Therefore, we must refresh our understanding continuously to gather useful experiences and keep them up-to-date.
6. Conclusion
The application of large-scale generative AI models has pushed forward the frontier of research in data storytelling tools. In this paper, we reflected the research progress through discussing the emerging and persistent human-AI collaboration patterns and their implications. We also identified the opportunities to advance this field. Due to space limit, this paper cannot discuss all exciting collaboration opportunities between humans and these AI models, such as applying video generation models (e.g., Sora (OpenAI, 2024)) for cinematic effects (Xu et al., 2023, 2022) or animated glyphs (Ying et al., 2022) and leveraging multi-agent systems (Shen et al., 2024a). Additionally, more roles of human and AI are worth exploration (Zhou et al., 2024b). We hope this paper can ignite future discussion about effective human-AI collaboration for data storytelling in the large-scale generative AI era.
Appendix A Visualizing the Trend of Data Storytelling Tools
Figure 2. The number of data storytelling tools between 2010 and 2024
[…]
This figure is a stacked bar chart to show the number of data storytelling tools between 2010 and 2024. The numbers across years are: 2 in 2010, 2 in 2013, 2 in 2014, 4 in 2016, 1 in 2017, 4 in 2018, 8 in 2019, 7 in 2020, 15 in 2021, 11 in 2022, 11 in 2023, and 20 in 2024. Among the 11 tools in 2023, 4 were collected by Li et al. and 7 were collected by us. The 20 tools in 2024 were fully collected by us.
Appendix B Visualizing the Relative Frequencies of Human-AI Collaboration Patterns
(a) Analysis
(b) Planning
(c) Implementation
(d) Communication
Figure 3. The human-AI collaboration pattern frequencies of earlier tools.
[…]
This figure shows the frequencies of collaboration patterns of earlier tools. The four subfigures show the pattern frequencies in four stages. In the analysis stage, the pattern of human-creator is 38%, human-creator and AI-assistant is 19%, AI-creator and human-assistant is 6%, AI-creator and human-optimizer is 25%, and AI-creator only is 12%. In the planning stage, the pattern of human-creator only is 52%, human-creator and AI-assistant is 10%, human-creator and AI-optimizer is 7%, AI-creator and human-optimizer is 28%, and AI-creator only is 3%. In the implementation stage, the pattern of human-creator only is 22%, human-creator and AI-assistant is 21%, AI-creator and human-assistant is 5%, AI-creator and human-optimizer is 34%, AI-assistant is 2%, AI-reviewer only is 3%, and AI-creator only is 12%. In the communication stage, the pattern of human-creator only is 60% and human-creator and AI-assistant is 40%. The other patterns that are not mentioned above have not been implemented in earlier tools.
(a) Analysis
(b) Planning
(c) Implementation
(d) Communication
Figure 4. The human-AI collaboration pattern frequencies of the latest tools.
[…]
This figure shows the frequencies of collaboration patterns of the latest tools. The four subfigures show the pattern frequencies in four stages. In the analysis stage, the pattern of human-creator is 14%, human-creator and AI-assistant is 21%, AI-creator and human-reviewer is 7%, AI-creator and human-optimizer is 21%, and AI-creator only is 36%. In the planning stage, the pattern of human-creator only is 35%, human-creator and AI-assistant is 12%, AI-creator and human-optimizer is 18%, AI-creator and human-reviewer is 6%, and AI-creator only is 29%. In the implementation stage, the pattern of human-creator only is 8%, human-creator and AI-assistant is 31%, AI-creator and human-optimizer is 27%, and AI-creator only is 35%. In the communication stage, the pattern of human-creator and AI-assistant is 33% and AI-creator only is 67%. The other patterns that are not mentioned above have not been implemented in the latest tools.